OpenStudy (anonymous):

How do I convert the polar equation r=1-cos theta into a rectangular equation?

OpenStudy (anonymous):

This question is on an exam review written this way? So I think it is correct.

OpenStudy (anonymous):

I can do coordinate conversions too. Thanks for trying.

OpenStudy (anonymous):

Anyone else?

OpenStudy (jdoe0001):

lemme check a bit

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Still thinking?

OpenStudy (anonymous):

I am still stuck

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

gotten anywhere?

OpenStudy (anonymous):

No. I am working on a different one because I don't know what to do. Still hoping for help!

OpenStudy (anonymous):

Any ideas?

OpenStudy (jdoe0001):

I'll show thus far what I'm checking :)

OpenStudy (jdoe0001):

$$ r=1-cos(\theta) (\times r)\\ r^2=r-rcos(\theta)\\ \text{bear in mind } rcos(\theta)=x\\ r^2=r-x\\ \text{now if you keep in mind that } \boxed{r^2=x^2+y^2} \ \&\& \ \boxed{r=\sqrt{x^2+y^2}}\\ $$

OpenStudy (jdoe0001):

from there, I gather you'd just need to solve for "y" or f(x)

OpenStudy (jdoe0001):

looks a bit convoluted, but is thus far what works :|

OpenStudy (anonymous):

That method will give you a fair bit of algebra.

OpenStudy (anonymous):

Is there a better method?